OpenStudy (anonymous):

How many gallons of 7% paint be mixed with 1 gallon of 18.2% paint to obtain 8.6% paint

OpenStudy (anonymous):

7*x+18.2*1=8.6(1+x) 18.2-8.6=1.6x x=9.6/1.6=6

OpenStudy (anonymous):

Where are you getting the (1+x) from that equation?

OpenStudy (anonymous):

new solution has the quantity of x gallon of 7% and 1 gallon of 18.2% so total (1+x) gallon.

OpenStudy (anonymous):

if u got it then let me know please..

OpenStudy (anonymous):

yea i got those kind of mixture problems thanks